OpenStudy (prettygirlnikki):

The 9th and 14th terms of an arithmetic sequence are -67 and -97,repectively. What is the 30th term?

OpenStudy (amistre64):

we can construct and equivalent sequence where -67 is the 0th term and -97 is the (14-9)th term to determine the (30-9)th term all the same

OpenStudy (amistre64):

bn = -67 +nd b0 = -67 + 0(d) = -67 b5 = -67 +5d = -97 d=(-97+67)/5 therefore: b21 = -67 +21d